OpenStudy (anonymous):

if f(x)= -2x-3 then what is the value of 5[f(d)] ?

OpenStudy (anonymous):

5[f(d)] =-10d-15

jimthompson5910 (jim_thompson5910):

f(x)= -2x-3 f(d)= -2d-3 5*[f(d)]= 5(-2d-3) 5*[f(d)]= -10d-15

OpenStudy (anonymous):

5 (-2d-3) = -10d -15

OpenStudy (anonymous):

thanks.